40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 96 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  12256   Wed Jul 6 20:51:00 2016 KojiUpdateGeneralSeismometer Enclosure Development | SURF 2016

Circuit1: It is nice to receive the voltage across the transimpedance resistor with a high impedance buffer (or amplifier), as close to the resister as possible. This amplifier needs to have low numbers for input bias current, input offset current, and input current noise. These current noise becomes the noise of the temperature reading. On the top of that, the input voltage noise of the buffer will be added to the output. The typical noise model can be found in http://www.analog.com/media/en/technical-documentation/application-notes/AN-940.pdf

The good candidates for the buffer is LT1128, ADA4004, OPA140, and LT1012. If the application is not too sensitive to the total noise, OPA604 is a good choise with easier handling.

Circuit2: With the same reason, AD741 is an old generic amp that is not a great choise for this purpose. The current noise is more significant because of the higher transimpedance here. The same noise model as above can be used to analyze the performance.

  12255   Wed Jul 6 19:36:45 2016 KojiUpdateGeneralSUS Vmon

I wanted to know what this Vmon exactly is. D010001 is telling us that the Vmon channels are HPFed with fc=30Hz (Attachment 1). Is this true?


I checked the quiscent noise spectrum of the ITMX UL coil output (C1:SUS-ITMX_ULCOIL_OUT) and the corresponding VMON (C1:SUS-ITMX_ULVmon). (Attachment 2 Ref curves). I did not find any good coherence. So the nominal quiscent Vmon output is carrying no useful information. 

Question: How much do we need to excite the coil output in order to see any meaningful signal?

As I excite the ITMX UL coil (C1:SUS-ITMX_ULCOIL_EXC) with uniform noise of 100-300 counts below 0.3Hz, I eventually could see the increase of the power spectrum and the coherence (Attachment 2). Below 0.1 Hz the coherence was ~1 and the transfer function was measured to be -75dB and flat. But wait, why is the transfer function flat?

In fact, if I inject broadband noise to the coil, I could increase the coil output and Vmon at the same time without gaining the coherence. (Attachment 3). After some more investigation, I suspect that this HPF is diabled (= bypassed) and aliasing of the high freq signal is causing the noise in Vmon.

In order to check this hypothesis, we need to visit the board.

Attachment 1: HPF.png
HPF.png
Attachment 2: 160706_ITMX_VMON2.pdf
160706_ITMX_VMON2.pdf
Attachment 3: 160706_ITMX_VMON1.pdf
160706_ITMX_VMON1.pdf
  12254   Wed Jul 6 17:17:22 2016 PrafulUpdateComputer Scripts / ProgramsNew Tabs and Working Summary Pages

The main C1 summary pages are back online now thanks to Max and Duncan, with a gap in pages from June 8th to July 4th. Also, I've added my new VMon and Sensors tabs to the SUS parent tab on the main pages. These new tabs are now up and running on the July 7th summary page.

Here's a link to the main nodus pages with the new tabs: https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20160707/sus/vmon/

And another to my ldas page with the tabs implemented: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1150848017-1150848317/sus/vmon/

Let me know if you have any suggestions or see anything wrong with these additions, I'm still working on getting the scales to be right for all graphs.

  12253   Wed Jul 6 16:40:09 2016 AakashUpdateGeneralSeismometer Enclosure Development | SURF 2016

I am using AD592CNZ temperature transducer ICs for measuring temperature inside as well as outside the enclosure. It is a  current output IC which outputs current proportional to temperature. As mentioned in the data sheet of AD592, I am using the following two schematics:

 

Though I still need to calibrate these temperature transducers, I did some measurements. I have temperature readings, and now my goal in few days is to find a transfer function of temperature fluctuations inside the enclosure to outside the enclosure.

 

About data acquisition:

We have re-configured the raspberry pi(B8:27:EB:70:D0:D8) on martian network. It's new ip address is 192.168.113.107(domenica.martian). Also, we have added the Acromag Busworks card(00:01:C3:00:9F:C8) on the martian network and its ip address is 192.168.113.237(acroey.martian).   

  12252   Wed Jul 6 11:02:41 2016 PrafulUpdateComputer Scripts / ProgramsVMon Tab on Summary Pages

I've added a new tab for VMon under the SUS parent tab. I'm still working out the scale and units, but let me know if you think this is a useful addition. Here's a link to my summary page that has this tab: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1151193617-1151193917/sus/vmon/


I'll have another tab with VMon BLRMS up soon.

Also, the main summary pages should be back online soon after Max fixed a bug. I'll try to add the SUS/VMon tab to the main pages as well.

  12251   Wed Jul 6 10:58:29 2016 VarunUpdateCDSDAFI review

The DAFI block was reviewed by Rana yesterday. The following changes/improvements were suggested: (Updated on 20th July 2016 with tasks taat remain in red)

1) include all the various channels like PEM, LSC, ASC, SUS, SEI, etc. as the inputs. Currently the inputs are only the LSC.
2) include all the control signals.
3) create a very detailed diagram of the entire signal flow and plan tasks accordingly.
4) Enable cascading of various DSP processes.
5) Adjusting the gain of the AGC such that the amplitude of the output signal comes to about half the peak amplitude offered by the ADC. This will help taking advantage of the entire dynamic range of the ADC.
6) change the enable button styles from a text input based controller to a button controller.
7) Currently, disabling a particular signal terminates the signal. Instead, it should turn into a unity gain block on disabling.
8) Check if the Fibox does AC coupling or not. If not, add an AC coupling arrangement in the DAFI.
9) Check the nature of the ADC1 and ADC2 inputs to the DAFI. I checked them yesterday, and they are channels 25 and 26 of ADC0, which are empty.

  12250   Wed Jul 6 10:58:07 2016 SteveUpdateVACvent 78 valve configuration

Vacuum Status: Chamber Open

All chamber annuloses are vented.  Vac Monitor screen is not communicating with gauges. The valve position indicator are working.

RGA is pumped by Maglev through VM2

 

Attachment 1: CamberOpen.png
CamberOpen.png
  12249   Wed Jul 6 09:34:26 2016 SteveUpdateGeneralPRM sensing voltage LR

Yesterday Q noticed that PRM_sensor_LR was 0.098V  This actually went to ~ zero on 7-3

 

Attachment 1: PRM_LR.png
PRM_LR.png
  12248   Wed Jul 6 08:34:00 2016 SteveUpdateVACRGA scan at day 639

The last RGA scan of this pumpdown 78

Pressure plot of 640 days long pd 78

CC1 cold cathode gauge was jump started with an accidental pressure glitch, that you can see on P1 plot

 

Attachment 1: RGAscan639d.png
RGAscan639d.png
Attachment 2: pd78d640.png
pd78d640.png
  12247   Tue Jul 5 23:38:42 2016 gautamUpdateGeneralVent progress - ETMX SUS Coil driver electronics investigation

With Koji's help, I've hacked together an arrangement that will allow us to monitor the output of the coil driver to the UL coil. 

The arrangement consists of a short custom ribbon cable with female DB25 connectors on both ends - the particular wire sending the signal to the UL coil has a 100 ohm resistor wired in series, because the coil has resistance ~20ohm, and the output of the coil driver board has a series 200(?) ohm resistor, so by directly monitoring the voltage at this point, we may not see a glitch as it may register too small. Tangentially related: the schematic of the coil driver board suggests that the buffered output monitor has a gain of 0.5. 

To monitor the voltage, I use the board to which the 4 Oplev signals are currently hooked up. Channel 7 on this particular board (corresponding to ADC channel 30 on c1scx) was conveniently wired up for some prior test, so I used this channel. Then, I modified the C1SCX model to add a testpoint to monitor the output of this ADC. Then, I turned OFF the input on the coil output filter for the UL Coil (i.e. C1:SUS-ETMX_ULCOIL_SW1) so that we can send a known, controlled signal to the UL Coil by means of awggui. Next, I added an excitation at 5 Hz, amplitude 20 counts (as the signal to the coil under normal conditions was approximately of this amplitude) to the excitation channel of the same filter module, which is the state I am leaving the setup in for the night. I have confirmed that I see this 5Hz oscillation on the monitor channel I set up. Oddly, the 0 crossings of the oscillations happen at approximately -1000 counts and not at 0 counts. I wonder where this offset is coming from? The two points I am monitoring the voltage across is shown in the attached photograph - the black clip is connected to the lead carrying the return signal from the coil.

I also wanted to set up a math block in the model itself that monitors, in addition to the raw ADC channel, a copy from which the known applied signal has been cancelled, as presumably a glitch would be more obvious in such a record. However, I was unable to access the excitation channel to the ULCOIL filter from within the SCX model. So I am just recording the raw output for tonight...

Attachment 1: image.jpeg
image.jpeg
  12246   Tue Jul 5 21:51:35 2016 ericqUpdateGeneralVent progress

One glitch was seen to occur without a change in the output voltage monitors in ELOG 11744

  12245   Tue Jul 5 19:10:41 2016 ranaUpdateGeneralVent progress

I'd suggest clamping and moving it to the flow bench so you can inspect with a bright light. Then remove the wire and inspect the standoff, but hurry up with getting it in the soak bath so you can start on the cleaning of the other ones.

I wonder if we're really sure that its a mechanical problem with ETMX.

Gautam tells me that the local damping was always ON when looking for the jumps. This means that the coil driver was still hooked up and we can't rule out glitches in the DAC or the coil driver.

The UL OSEM shows the biggest movement (10 microns). The LR shows the second most (6-7 microns). The others are 2x less. So its consistent with a voltage change on UL,

Is this consistent with a slip in one of the wire standoffs? I think no.

 

Attachment 1: jumpEvent20151105.png
jumpEvent20151105.png
  12244   Tue Jul 5 18:44:39 2016 PrafulUpdateComputer Scripts / ProgramsWorking 40m Summary Pages

After hardware errors prevented me from using optimus, I switched my generation of summary pages back to the clusters. A day's worth of data is still too much to process using one computer, but I have successfully made summary pages for a timescales of a couple of hours on this site: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/

 

Currently, I'm working on learning the current plot-generation code so that it can eventually be modified to include an interactive component (e.g., hovering over a point on a timeseries would display the GPS time). Also, the 40m summary pages have been down for the past 3 weeks but should be up and working soon as the clusters are now alive.

  12243   Tue Jul 5 17:14:02 2016 ericqUpdateGeneralVent progress

Rough summary of today's progress:

  • IMC locked, reasonably aligned. (Only have MC2 trans reference, no WFS)
    • FSS Slow lock threshold got lowered to 1k MC2 trans
    • Autolocker doesn't really work, though I made some modifcations to it that reflect the manual locking that works
  • TTs, PRMI optics moved around until AS and REFL spots looked good. 
    • Small flashes visible in TRY
    • ND filter on Y transmon QPD removed, since we have so little light
  • Y arm was not locked, but this was deemed suitable to proceed
  • Heavy doors taken off ETMX and ITMX chambers, light doors put on.
  • Went into ETMX chamber, poked around, took a bunch of pictures (Gautam is uploading has uploaded to picasa (foteee account), the two best ones of the stand-offs on either side are attached, although I guess we can't really conclude anything from these except that on the side with the OSEM, there seems to be a few extra stray bits of epoxy residue...)
  • Removed ETMX side OSEM to get a picture of that side of optic (it is currently off, ETMX watchdog shutdown for now...)
  • Checked that table is level with clean spirit level (it was to within the resolution of the spirit level which is quoted as 0.42mm/m)

I didn't really see anything out of the ordinary on the ETMX suspension. Earthquake stops had clearance, OSEMS were secure, no visible glue degredation on face magnets. Inspection with green LED flashlight didn't reveal any obscene dirtieness on either face, just a few particles here and there. The top of the opic barrel unsurprisingly has a good amount of particulate. The wire grooves are way too small to resolve anything at this point, other than that they exist.

The suspension footprint is already marked, tomorrow we can move the suspension closer to the door to get an even closer look at it, before removing it from the chamber. 

Attachment 1: noOsemSide.JPG
noOsemSide.JPG
Attachment 2: OSEMside.JPG
OSEMside.JPG
  12242   Tue Jul 5 14:12:56 2016 varunUpdateElectronicsAntialiasing Filter Update

I am trying to design an antialiasing filter, which also has two switchable whitening stages. I have designed a first version of a PCB for this.

The board takes differential input through PCB mountable BNCs. It consists of an instrumentaiton amplifier made using quad opamp ADA4004, followed by two whitening blocks, also made using ADA4004, which can be bypassed if needed, depending upon a control input. The mux used for this purpose is Maxim MAX4158EUA. These two whitening blocks are followed by 2 the LPF stages. A third LPF stage could be added if needed. These use AD829 opamps. After the LPFs are two amplifiers for giving a differential output through two output BNCs. The schematic is shown in attachment 1: "AA.pdf". The top layers of the layout are shown in attachment 2 (AAtop.pdf), the bottom layers in attachment 3 (AAbottom.pdf), and the entire layout in attachment 4 (AAbrd.pdf). 

The board has 6 layers (in the order from top to bottom):

1) Top signal layer; 

2) Internal plane 1 (GND),

3) Internal plane 2 (+15V),

4) Internal plane 3 (-15V),

5) Internal plane 4 (GND),

6) Bottom signal layer. 

Power: +15, -15 and GND is given through a 4 pin header connector. 

The dimensions of the board are 1550 mil \times 6115 mil (38.1mm\times155.3mm) and the overall dimensions including the protruding BNC edges are 1550 mil \times 7675 mil (38.1mm\times194.9mm)

I would like to have inputs on the layout telling me if any component/trace needs to be changed/better placed, any other things about the board need to be changed, etc.

 

P.S.: I have also added a zipped folder "AA.zip" containing the schematic and board files, as well as the above pdfs.

Attachment 1: AA.pdf
AA.pdf
Attachment 2: AAtop.pdf
AAtop.pdf
Attachment 3: AAbottom.pdf
AAbottom.pdf
Attachment 4: AAbrd.pdf
AAbrd.pdf
Attachment 5: AA.zip
  12241   Tue Jul 5 09:25:31 2016 SteveUpdateSUS how to remove epoxy from optics

Glass soaking dish with teflon guides.

Quote:

Proposed Acetone soak dish for SOS epoxy softening.

It has good acces through 5" top ID. The set up is stable and teflon lined.

Materials: glass jar with SS cover, teflon bricks, 0.008" teflon wrapped "high density Drever bricks" and aluminum

Drever brick: I beleive it is a Tungsten alloy. We used it as vac-bat savor at the coffe can. It has high density, heavy and hard, it was never identified.

I will soak one brick  to see if it has any reaction ability with acetone.

 

NO means that only Glass and Teflon can be used for this fixture in  Acetone. We can not take a chance on the coating!

I guess the small surface area Aluminum dumbbell, guide rod and-or wire standoff, magnet and epoxy does not degrade the acetone such way that it effects our coating.

Not ot mention, that only the very edge of the coating would in this solution.

              

 

 

Attachment 1: SOSsoakingDish.jpg
SOSsoakingDish.jpg
Attachment 2: SOSsoakingD.jpg
SOSsoakingD.jpg
  12240   Mon Jul 4 10:30:37 2016 ericqUpdateGeneralPreliminary vent plans

Here are some plans / rough procedures for this week's vent. It is unlikely that I have though of everything, but this should be a reasonable starting point.

The mode cleaner still hasn't been locked in air, we may not want to touch the Y arm optics until we are able to lock to the Y arm and dither align, so we are sure to keep the input pointing from drifting away too much.


Primary objectives:

  • Re-suspend ETMX
  • First contact of all arm cavity optics

Secondary objectives:

  • Install new gauges
  • Replace 40mm baffles with 50mm baffles
  • Check cleanliness of inner viewport surfaces

ETMX project

  • Open ETMX chamber
  • Take all manner of photos of ETMX suspension in-situ
    • If some kind of obvious issue is evident, fix it, proceed accordingly
  • Mark suspension position
  • Move suspension to edge of the table, more pictures + inspection
  • Move suspension to flow bench, remove optic
  • Transport optic to clean room
  • Acetone soaking / standoff removal
  • Re-glue side magnet
  • Re-glue guide rod + standoff
  • OSEM transplant from old to new suspension
  • Suspend, following SOS suspension procedure
  • Drive optic around, see if jumps are evident
  • Clean with first contact
  • Reinstall optic, align, etc.

Optic cleaning

For $optic in [ITMX, ITMY, ETMY]:

  • Open chamber
  • Take many, many pictures
  • Mark suspension position
  • move suspension to edge, take pictures of HR surfaces
  • Mark OSEM orientation, remove oems
  • Clean AR and HR surfaces with first contact
  • Reinstall OSEMS at proper position, rotate to minimize bounce/roll coupling
  • Reinstall optic, align, etc.
  12239   Fri Jul 1 17:51:28 2016 PrafulSummaryElectronicsReplacing DIMM on Optimus

There has been an ongoing memory error in optimus with the following messages:

controls@optimus|~ >
Message from syslogd@optimus at Jun 30 14:57:48 ...
 kernel:[1292439.705127] [Hardware Error]: Corrected error, no action required.

Message from syslogd@optimus at Jun 30 14:57:48 ...
 kernel:[1292439.705174] [Hardware Error]: CPU:24 (10:4:2) MC4_STATUS[Over|CE|MiscV|-|AddrV|CECC]: 0xdc04410032080a13

Message from syslogd@optimus at Jun 30 14:57:48 ...
 kernel:[1292439.705237] [Hardware Error]: MC4_ADDR: 0x0000001ad2bd06d0

Message from syslogd@optimus at Jun 30 14:57:48 ...
 kernel:[1292439.705264] [Hardware Error]: MC4 Error (node 6): DRAM ECC error detected on the NB.

Message from syslogd@optimus at Jun 30 14:57:48 ...
 kernel:[1292439.705323] [Hardware Error]: cache level: L3/GEN, mem/io: MEM, mem-tx: RD, part-proc: RES (no timeout)

Optimus is a Sun Fire X4600 M2 Split-Plane server. Based on this message, the issue seems to be in memory controller (MC) 6, chip set row (csrow) 7, channel 0. I got this same result again after installing edac-utils and running edac-util -v, which gave me:

mc6: csrow7: mc#6csrow#7channel#0: 287 Corrected Errors 

and said that all other DIMMs were working fine with 0 errors. Each MC has 4 csrows numbered 4-7. I shut off optimus and checked inside and found that it consists of 8 CPU slots lined up horizontally, each with 4 DIMMs stacked vertically and 4 empty DIMM slots beneath. I'm thinking that each of the 8 CPU slots has its own memory controller (0-7) and that the csrow corresponds to the position in the vertical stack, with csrow 7 being the topmost DIMM in the stack. This would mean that MC 6, csrow 7 would be the 7th memory controller, topmost DIMM. The channel would then correspond to which one of the DIMMs in the pair is faulty although if the DIMM was replaced, both channels 0 and 1 would be switched out. Here are some sources that I used:

http://docs.oracle.com/cd/E19121-01/sf.x4600/819-4342-18/html/z40007f01291423.html#i1287456

https://siliconmechanics.zendesk.com/hc/en-us/articles/208891966-Identify-Bad-DIMM-from-EDAC

http://martinstumpf.com/how-to-diagnose-memory-errors-on-amd-x86_64-using-edac/

I'll find the exact part needed to replace soon.

  12238   Fri Jul 1 15:48:10 2016 SteveUpdateVACvent 78 at 710 Torr

I just disconnected the 6th instrument grade air cylinder from the vacuum envelope at 720 Torr. Now it will reach equilibrium through a filter as it sucks in lab air.

This is the sure way not to over pressurize the chamber.

 

Attachment 1: pd78at710T.png
pd78at710T.png
  12237   Fri Jul 1 09:20:44 2016 SteveUpdateVACvent 78

  VENT OBJECTIVES:

1, Fix ETMX sus "jump issue"

2, First Contact clean the arms

3, Install new spare cold cathode and convectron gauges: InstruTech-Hornet

4, Install 50 mm apeture beam baffles

5, Check and clean optical quality viewport from inside

 

The following bullets were executed:

  • Closed shutter of PSL-IR and green shutters at the ends 
  • Checked jam nuts are protecting bellows 
  • Check crane functionality & cleanliness last week

  • Particle count must be under 10,000 counts / cf min for 0.5 micron 
  • Checked all metal window covers are on. 
  • Check 5 cylinders (24 cft size) of instrument grade air, called Alfa Gas 1 in stock.
  • Took pictures of medm screens: sus-sum, aligned oplevs pos, alignment values and vac configuration
  • Turned Oplev servos off
  • Closed V1 and VM1, opened VM2
  • Opened VV1 manual valve to Nitrogen cylinder and let the P1 to rise to 25 Torr
  • Switched over to Instrument Grade Air Cylinder and continoed the vent with 8-10 PSI reading at the pressure regulator

We are venting the 40m IFO

     

Quote:

Steve has ordered some teflon parts to take the place of the metal parts in his acetone-soaking jig. They should arrive tomorrow. 

So, we will be begin the venting process tomorrow. Doors to come off on Tuesday.

 

Attachment 1: vent78.png
vent78.png
Attachment 2: vent78at250T.png
vent78at250T.png
  12236   Fri Jul 1 01:52:54 2016 AakashUpdateGeneralSeismometer Enclosure Development | SURF 2016

I have transferred most of the temperature measurement stuff from the front area to seismometer at the end of Y-arm.  While arranging the components I have taken all care that they will not interfere with existing system. Also, I have temporarily taken a monitor from the front area to the area near same seismometer as I couldn't talk to Rpi via ssh. For next twelve hours, I am now recording temperature inside as well as outside the seismometer enclosure. Some temperature sensors are inside the enclosure while some are outside the seismometer enclosure.

 

  12235   Thu Jun 30 21:24:36 2016 ericqUpdateGeneralVent Prep

Steve has ordered some teflon parts to take the place of the metal parts in his acetone-soaking jig. They should arrive tomorrow. 

So, we will be begin the venting process tomorrow. Doors to come off on Tuesday.

  12234   Thu Jun 30 16:21:32 2016 gautamUpdateCOCSideband HOMs resonating in arms

[EricQ, gautam]

Last night, we set about trying to see if we could measure and verify the predictions of the simulations, and if there are indeed HOM sidebands co-resonating with the carrier. Koji pointed out that if we clip the transmitted beam from the arm incident on a PD, then the power of the higher order HG modes no longer integrate to 0 (i.e. the orthogonality is broken), and so if there are indeed some co-resonating modes, we should be able to see the beat between them on a spectrum analyzer. The procedure we followed was:

  1. Choose a suitable PD to measure the beat. We chose to use the Thorlabs PDA10CF because it has ~150MHz bandwidth, and also the responsivity is reasonable at 1064nm.
  2. We started our measurements at the Y-end. There was a sufficiently fast lens in the beam path between the transmon QPD and the high gain PD at the Y end, so we went ahead and simply switched out the high gain thorlabs PDA520 for the PDA10CF. To power the PDA10CF, we borrowed the power cable from the green REFL PD temporarily.
  3. We maximized the DC power of the photodiode signal using an oscilloscope. Then to introduce the above-mentioned clipping and orthogonality-breaking, we misaligned the beam on the PD until the DC power was ~2/3 the maximum value. 
  4. We then hooked up the PD output to the Agilent network analizyer (with a DC block).
  5. We measured the spectrum of the PD signal around 11.066MHz (with 100kHz span) and higher harmonics up to 55MHz and used a narrow bandwidth (100Hz) and long integration time (64 averages) to see if we could find any peaks. More details in the results section.
  6. Having satisfied ourselves with the Y-end measurements, we 
  • restored the power cable to the green beat PD
  • re-installed the thorlabs PDA520 
  • verified that both IR and green could be locked to the arm

We then repeated the above steps at the X-end (but here, an additional lens had to be installed to focus the IR beam onto the PDA10CF - there was, however, sufficient space on the table so we didn't need to remove the PDA520 for this measurement).


Results:

Y-end: DC power on the photodiode at optimal alignment ~ 200mV => spectra taken by deliberately misaligning the beam incident on the PD till the DC power was ~120mV (see remarks about these values).

RF sideband (Y-arm) Peak height (uV) Beat power (nW) RF sideband (X-arm) Peak height (uV) Beat Power (nW)
11 1.55 0.52 11 1.2 0.4
22 10.6 3.53 22 none seen N.A.
33 none seen N.A. 33 none seen N.A.
44 22.0 7.33 44 7 2.33
55 8.6 2.97 55 5 1.67

I converted the peak heights seen on the spectrum analyzer in volts to power by dividing by transimpedance (=5*10^3 V/A into a 50ohm load) * responsivity at 1064nm (~0.6A/W for PDA10CF).


Remarks:

  1. This effect flagged by the simulations seems to be real. Unfortunately I can't get a more quantitative picture because we can't quantify the mode-overlap between the carrier 00 mode and any higher order mode on the beat PD (as we know nothing about the profile of these modes), but the simulations did suggets that the 2nd order 22MHz and 4th order 44MHz HOMs are the ones closest to the carrier 00 resonance (see Attachments #2 and #3), which is kind of borne out by these results. 
  2. I disbelieve the conversions into power that I have done above, but have just put them in for now, because a DC power of 200mW at the Y-end suggests that there is >160uW of light transmitted from the arm, which is at least twice what we expect from a simple FP cavity calculation with the best-known parameters. If I've missed out something obvious in doing this conversion, please let me know! 
  3. For the Y-arm, the region around 55MHz had a peak (presumably from the sideband HOM beating with the carrier) but also a bunch of other weird sub-structures. I'm attaching a photo of the analyzer screen. Not sure what to make of this...
Attachment 1: image.jpeg
image.jpeg
Attachment 2: C1_HOMcurves_Y.pdf
C1_HOMcurves_Y.pdf
Attachment 3: C1_HOMcurves_X.pdf
C1_HOMcurves_X.pdf
  12233   Thu Jun 30 16:11:57 2016 ericqUpdateGeneralVent Prep

I have updated the vent prep checklist on the wiki. Gautam and I did the following things from it:

  • Center all oplevs, transmon QPDs 
    • ETMX oplev has not been centered, since it's moving around so much, and we're going to immediately move the suspension anyways.
  • Align the arm cavities for IR and align the green lasers to the arms. 
    • AUX X Green was aligned while the X arm was well aligned. Soon thereafter, ETMX wandered away, but the green will remain a good reference
  • Update the SUS Dritmon values 
  • Reconcile all SDF differences 
  • Reduce input power to no more than 100mW by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.) 

  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
    • I don't think we've vented since the most recent slew of changes to the IMC servo, so its not surprising that the current low power scripts don't work. I'm working on locking the IMC, but this does not prevent us from initiating the vent tomorrow.

 

The following bullets have not yet been executed:

  • Close shutter of PSL-IR and green shutters at the ends 
  • Make sure the jam nuts are protecting bellows 
  • Check crane functionality & cleanliness 

  • Turn off HV into vacuum: OMC is not wired this time 
  • Particle count must be under 10,000 counts / cf min for 0.5 micron 
  • Check all metal window covers are on. 
  • Check 5 cylinders (24 cft size) of instrument grade air, called Alfa Gas 1 in stock.
  12232   Thu Jun 30 14:31:02 2016 ChemistryUpdateSUSNO

  12231   Thu Jun 30 10:03:16 2016 SteveUpdateSUS how to remove epoxy from optics

Proposed Acetone soak dish for SOS epoxy softening.

It has good acces through 5" top ID. The set up is stable and teflon lined.

Materials: glass jar with SS cover, teflon bricks, 0.008" teflon wrapped "high density Drever bricks" and aluminum

Drever brick: I beleive it is a Tungsten alloy. We used it as vac-bat savor at the coffe can. It has high density, heavy and hard, it was never identified.

I will soak one brick  to see if it has any reaction ability with acetone.

 

NO means that only Glass and Teflon can be used for this fixture in  Acetone. We can not take a chance on the coating!

I guess the small surface area Aluminum dumbbell, guide rod and-or wire standoff, magnet and epoxy does not degrade the acetone such way that it effects our coating.

Not ot mention, that only the very edge of the coating would in this solution.

              

 

Attachment 1: SOSsoakDish.jpg
SOSsoakDish.jpg
Attachment 2: SOSsoakDish_tv.jpg
SOSsoakDish_tv.jpg
  12230   Wed Jun 29 16:05:47 2016 ericqUpdateSUSETMX suspension readiness

I've gone through the SOS suspension document (E970037) and some old elogs to get an idea of all the accesories we need for the process of suspending, aside from the tower itself, which Steve has already put together. Gautam and I have laid our eyes upon most of the critical pieces. Some other objects are unknown, and perhaps not strictly neccesary. 

Confirmed to exist:

  • Guide rod gluing fixture
  • 3 axis micrometer + microscope doohicky
  • PZT buzzer (for fine standoff positioning)
  • Winch fixture (the plate maybe doesn't fit perfectly on the top of the SOS tower, but one bolt will probably work)
  • Spare guide rods and ruby standoffs
  • Magnet+dumbell gluing fixture (in case we need to glue a new magnet+dumbell pair)
  • Magnet to optic gluing fixture (including "pickle pickers"/grippers)
  • 0.017" suspension wire
  • HeNe laser + QPD + beam height target (confirmed working on scope)

In addition, I am told that we have a long ribbon cable that can run from the X end to the clean room to enable OSEM damping control while we do the pitch alignment.

Things mentioned in the procedure I have not found:

  • Something called an "SOS set screw tool" (D961412)
    • Not mentioned specifically what this is for
  • Edmund scientific pocket measuring microscope + bushing
    • Seems to be intended for centering OSEM holding plate on the magnet position. I found no mention of this use in 40m elogs. It can likely be done by eye
  • SOS cleaning bracket (D970181)
    • This is likely for the old liquinox cleaning procedure. We'll be using first contact.

Some other tasks and their status:

  • Check cleanroom table levelling:
    • Done via 12" spirit level, looks good
  • Glue removal strategy / fixture
    • Steve is working on this
  • Cut and apply viton tips to earthquake stops
    • Either gautam will do this, or we will use the current suspension's stops
  12229   Wed Jun 29 16:02:29 2016 SteveUpdateSUS ETMX beam baffle

In the attached photo from 2012, one can see the installed black glass baffle. According to the drawings (LIGO-DNNNXXX) this one has a clear aperture of 40 mm.

In (someplace ?) we have clean baffles with a 50 mm aperture which can be installed during this vent. In order to be more conservative, let us choose to swap these out for all 4 test masses during the upcoming vent using the green laser as an alignment guide, as Koji described at today's lunch meeting.

They are located at the top of E1 drawer cabinet

Attachment 1: IMG_1608.JPG
IMG_1608.JPG
  12228   Wed Jun 29 15:34:00 2016 SteveUpdatesafetySURF 2016 safety

Praful Vasiceddy received 40m specific basic safety training.

Quote:

Aakash Patil received 40m specific basic safety training.

Quote:
Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

Varun has received 40m specific basic safety training today.

 

 

  12227   Wed Jun 29 14:32:22 2016 SteveUpdateSUS how to remove epoxy from optics

 

Hi Steve -

I found the doc I was looking for:

https://dcc.ligo.org/LIGO-E1200821

Specifically, you might find guidance in Section 5 and the pictures at the end of the doc.  This should work for Vacseal as well.

Good luck - it will take some time (hours to day or 2)...
I'll be interested to know how it goes.

GariLynn helped us develop this procedure so you could also ask her to cast an eye over the setup if you are worried.

-Betsy

ps: there is no existing fixture to hold SOS optic while soaking it
 

  12226   Wed Jun 29 08:20:04 2016 SteveUpdateSUS First Contact shelf life
We'll follow LIGO policy:
Our policy is to use first contact within 1 year of purchase for use in the interferometers.  For inspection use I am comfortable with out-of-date use.

GaryLinn offered their indate First Contact for use.

 

  12225   Wed Jun 29 00:09:36 2016 AakashUpdateGeneralThings from past | SURF 2016

I have taken out the heaters and temperature sensors from the enclosure which was made by Megan last summer. Soon I will test and configure those heaters.

  12224   Tue Jun 28 22:54:43 2016 AakashUpdateGeneralSeismometer Enclosure Development | SURF 2016

The existing enclosure for seismometer at LIGO 40m lab is a cylindrical stainless steel can placed upside down over the seismometer. It has more empty space between the seismometer and the internal surface of enclosure which is not desirable(I'll quantitatively elaborate this statement once my temperature measuring setup is ready).

 

Stainless steel has a thermal conductivity in the range of 16.3 to 16.7 W/m/K and magnetic permeability 1.260e-6 H/m.Assuming an ambient temperature 298K, and the temperature inside the enclosure as 295K, as well as substituting all the values for dimesions and material properties of existing enclosure,
k=16.4 W/mK, μ=1.260e-6 H/m, L=2ft=0.6096m, b=r2 =0.5ft=0.1524m, thickness=5mm, a=r1 =0.1474m.
So by using the textbook relations(I have mentioned them in my report), the value of attenuation coefficient is 5.953584e-05 and the value of rate of heat transfer= 5.64913 kW. The attenuation coefficient value is quite better for steel but proper care needs to be taken to avoid heat transfer. For studying the variation of rate of heat transfer and attenuation with the thickness of enclosure material, I have plotted the following attached graphs for different materials which include hardened stainless steel, aluminium, pure iron and nanoperm-muMetal.

 

 

About Data Acquisation

I have already invested a lot of time to configure and use acromag busworks card over ethernet. So now I have made an arrangement to measure temperature by AD592CNZ temperature transducer IC. I would be using raspberry pi for acquiring data untill I figure out a way to use acromag busworks card for the same. This setup of acquiring logging temperature using raspberry pi is mostly ready except the calibration part.

  12223   Tue Jun 28 20:43:23 2016 KojiSummaryCOCFirst Contact cleaning practice

Made a dry run of the in-situ cleaning for a 3inch optic.

Attachment 1: The Al dummy mass is clamped in the suspension cage.
Attachment 2: The front surface was painted. The nominal brush with the FC bottle was used.
Attachment 3: Zoom in of the front surface.
Attachment 4: The back surface was painted.
Attachment 5: The back surface was peeled.
Attachment 6: The front surface was peeled too.
Attachment 7: The peeled layers.

Findings:

1. To paint a thick layer (particlarly on the rim) is the key to peel it nicely.

2. It was helpful for easier peeling to have mutiple peek tabs. Two tabs were sufficient for ~1" circle.

3. The nominal brush with the bottle was OK although one has to apply the liquid many times to cover such a large area. A larger brush may cause dripping.

4. The nominal brush was sufficiently long once the OSEMs are removed. In any case it is better to remove the OSEMs.

Attachment 1: IMG_20160628_170335196.jpg
IMG_20160628_170335196.jpg
Attachment 2: IMG_20160628_171547769.jpg
IMG_20160628_171547769.jpg
Attachment 3: IMG_20160628_171607802.jpg
IMG_20160628_171607802.jpg
Attachment 4: IMG_20160628_172328190.jpg
IMG_20160628_172328190.jpg
Attachment 5: IMG_20160628_174541960.jpg
IMG_20160628_174541960.jpg
Attachment 6: IMG_20160628_174556004.jpg
IMG_20160628_174556004.jpg
Attachment 7: IMG_20160628_174617198.jpg
IMG_20160628_174617198.jpg
  12222   Tue Jun 28 17:11:27 2016 PrafulUpdateGeneralEM172 Microphones

Found 60 EM172 microphones. Previous elog with details: 7777.

  12221   Tue Jun 28 16:10:49 2016 PrafulUpdateGeneralBluebird Microphones

Found 1 out of 2 bluebird microphones in the 40m.

  12220   Tue Jun 28 16:09:41 2016 PrafulUpdateGeneral40m Summary Pages

Set up gwsumm on optimus and generated summary pages from both L1 and C1 data. Still a few manual steps need to be taken during generation, not fully automated due to some network/username issues. nds2 now working from optimus after restarting nds2 server.

  12219   Tue Jun 28 16:06:09 2016 gautamUpdateCOCRC folding mirrors - further checks

Having investigated the mode-overlap as a function of RoC of the PRC and SRC folding mirrors, I've now been looking into possible stability issues, with the help of some code that EricQ wrote some time back for a similar investigation, but using Finesse to calculate the round trip Gouy phase and other relevant parameters for our current IFO configuration. 

To do so, I've been using:

  1. Most up to date arm length measurements of 37.81m for the Y arm and 37.79m for the X arm
  2. RoCs of all the mirrors from the phase map summary page
  3. Loss numbers from our November investigations

As a first check, I used flat folding mirrors to see what the HOM coupling structure into the IFO is like (the idea being then to track the positions of HOM resonances in terms of CARM offset as I sweep the RoC of the folding mirror). 

However, just working with the flat folding mirror configuration suggests that there are order 2 22MHz and order 4 44MHz HOM resonances that are really close to the carrier resonance (see attached plots). This seems to be originating from the fact that the Y-arm length is 37.81m (while the "ideal" length is 37.795m), and also the fact that the ETM RoCs are ~3m larger than the design specification of 57m. Interestingly, this problem isn't completely mitigated if we use the ideal arm lengths, although the order 2 resonances do move further away from the carrier resonance, but are still around a CARM offset of +/- 2nm. If we use the design RoC for the ETMs of 57m, then the HOM resonances move completely off the scale of these plots... 

Attachment 1: C1_HOMcurves_Y.pdf
C1_HOMcurves_Y.pdf
Attachment 2: C1_HOMcurves_DR.pdf
C1_HOMcurves_DR.pdf
  12218   Tue Jun 28 09:14:27 2016 SteveUpdateSUS ETMX optics preparation to be rehanged

Given the effect on the contrast defect, the consensus at the meeting Wednesday 6-22  was that we should continue to use the existing ETMX optic. 

 
Does this imply that we can simply:
  • Vent
  • Remove the current ETMX suspension, and optic from the suspension
  • Replace the standoffs (and guide rod if necessary), leaving the current magnets in place
  • Re-suspend on the current, existing suspension, with new wire clamps
That is, if we are keeping the current optic, do we gain anything by replacing the whole suspension? Is there any reason to remove and replace magnets? (The wire clamps do need to be replaced, as a groove forms when they are installed.)
 
This could allow us to vent and execute this whole thing sooner. 
 
-Eric Q.
...............................
Yes, I think this is a reasonable approach. Besides the wire standoffs and wireclamps, the only other problem might by a worn groove in the top piece of the suspension (not the clamp piece, but the piece that the clamp holds the wire up against).
 
How about use the current ETM, but put it into the new suspension body with new wire standoffs?
 
Quote:

D Location

Number on

Drawing

 

Component

Name

 

 Baked Clean

 Pieces Needed

 

Pieces

In Stock

(on tower )

Notes
         
31,20,19,13 viton tips   not cut yet baked material in stock
30 6-32x0.75" stops 4 (4)  
29 steel music wire 0.017" not baked on roll  needs good wipes
28  1/4 "washer 4 (4)  
27 lock washer 4 50 install
26 Ag plated 1/4-20x1.25 4 (4)  
25 Ag plated1/4-25x.75 & not plated 20 (20)  
24 SS 4-40x.5 2 (2)  
23 SS 4-40x.38 4 (4)  
22 spring plunger 4+2 (4+2)  
         
         
18 magnets, 1.9mm od, length 3.2 mm 5 ~30 not coated, rusty buy Ni coated ones for future use
17 guide rod, 0.635 mm od, 3.3 mm 3 6 Al
16 wire standoff, 1 mm od, 4.8 mm 2 2 Al and ruby (ruby groove not centered)
15 short OSEMs 5 6  
14 spare ETMX in 40m wiki 1 1  confirmed in cabinet
         
12 dumbbell standoff 5 6  
11 Al stiffening plate 1 (1)  
10 wire clamp B in sus block 2 (2)  
9 wire clamp A 1 (1)  
8-7 lower clamp for lifting optics 2 (2)  
6 upper clamp to hold down optic 1 (1)  
5-4 left-right side of tower  1 ea (1ea)  
3 tower base 1 (1)  
2 sus block 1 (1)  
1 lower and upper OSEM holders 1ea (1ea)  
         
48 sandind fixture for magnet&dumbbell      
45 magnet-dumbbell assemblly fixture 1 1  
43 guide rod gluing fixture     shipped from LOH .... D000335...needs cleaning
         
  First contact   3-15-2013 purchase check expiration date OK
  FCPEEK peeler ring disk for TM cleaning 10 front,10 back side  have sheets only 32&19mm ID punches ordered
  GordonBrush custom for LIGO optical cleaning ~5 1 (3/8wide nylonSS) more from Calum available
  EP30-2 epoxy have  have expiration date 9-24-2016

NOT finished, last edited 6-28

 

Attachment 1: nylonFCss.jpg
nylonFCss.jpg
  12217   Mon Jun 27 15:47:17 2016 SteveUpdateGeneralQPR clean room gloves

I got some QPR Nitrile gloves. They are LIGO approved.White nitrile gloves are naturally anti-static- 109 ohms

Their touch not as good as laytex  gloves but try to use them.

 

  12216   Mon Jun 27 15:26:03 2016 SteveOmnistructureALARMfire alarm test

The fire alarm came on around 15:05  for about 2-3 minutes. We all  left the lab and counted heads.  I called Paul Mackel x2646 (cell 626/ 890- 3259) at Fire Protection Services. He said that this alarm test was planned and we should of got an email notice. Perhaps I missed that notes.

Quote:

Fire alarm went off several minutes ago. Talked to security and they said there was no fire. It beeped twice again just now. No one has been working on the IFO today.

 

Attachment 1: fireAlarmTest.png
fireAlarmTest.png
  12215   Mon Jun 27 15:12:09 2016 varunUpdateCDSDAFI update: stereo output

Using an RC to BNC connector from the inner drawer, I have added a second output cable going from the output Fibox in the control room to the audio mixer.

Quote:

I have updated the DAFI with the following changes:

1) Separated both the channels of stereo output completely, as well as in the GUI.

2) Added text monitors for the inputs and outputs.

The stereo output is now ready except for a cable going from the second channel of the output fibox to the audio mixer.

Attached is the main DAF_OVERVIEW screen and its link button from the LSC screen labelled "DAFI"

Quote:

I wish to have stereo audio output for the DAF module. Hence, there needs to be a second output from the DAF. I added this second output to the model. Following are the details:

FiBox: It consists of two analog inputs which are digitized and multiplexed and transmitted optically. (only 1 fiber is needed due to multiplexing). Attachment 1 shows the fibox with its 2 analog inputs (one of which, is connected), and 1 fiber output. The output of the DAF goes to the FiBox. Until today, the Fibox recieved only 1 analog input. This analog signal comes from the DAC-8 (count starting from 0), which is located at "CH 1 OUT" SMA output in the "MONITORS" bin on the racks (attachment 2).

I have added another output channel to the DAF model both in software and in hardware. The DAF now also uses DAC-9 analog output which goes to the second analog input of the FiBox. The DAC-9 output is located at "CH 2 OUT" SMA output in the "MONITORS" bin on the racks (attachment 4).

After making the changes, the Fibox is shown in attacment 3.

Testing: The LSC input on passing through the DAF block is given through two different DAC outputs, to the same Fibox channel (one after the other), and the output is heard. More concrete testing will be done tomorrow. It will be as follows:

1) Currently, I need to search for a suitable cable that would connect the second channel of the output fibox to the audio mixer. After doing this, end to end testing of both channels will be done.

2) I could not access the AWG, probably because the DAQ was offline today afternoon. Using a signal from the AWG will give a more concrete testing of the stereo output.

3) After this, I will separate the two channels of the stereo completely (currectly they are seperated only at the DAF output stage)

4) I also will edit the medm gui appropriately.

 

Quote:

I have added Enable buttons for each of the DSP blocks, and labels for the matrix elements. The input matrix takes inputs from each of the 4 channels: ADC1, ADC2, LSC and EXC, and routes them to the audio processing blocks (attachment 2). The output matrix (attachment 3) takes the outputs of the various DSP blocks and routes them to the output and then to the speakers. 

 

 

 

Attachment 1: IMG_20160627_151753247.jpg
IMG_20160627_151753247.jpg
  12214   Sun Jun 26 15:27:28 2016 ranaFrogsIOOPMC /MC lopced

Found PMC unlocked for many hours so I relocked it. IMC relocked by itself, but the input switch seems to be flickering to fast. Also the Keep Alive bit is not flashing. no

  12213   Thu Jun 23 17:24:32 2016 ericqUpdateGeneralTweaks

I spent some time this afternoon reviving some of my CESAR/ESCOBAR shenanigans on the Y arm. I found it neccesary to adjust a few things.

  • PMC realigned 
  • ETMY oplev centered
  • Y End green realigned
  • PSL/Y Green beat realigned

Afterwards, ALSY noise levels were good. 

  12212   Wed Jun 22 14:03:42 2016 VarunUpdateElectronicsAnti-Aliasing Filter circuit schematic

I found an anti-aliasing circuit on the 40m wiki. It consists of A differential LPF made using THS4131 low noise differential op-amp (one of the main applications of which is preprocessing before the ADC), and a notch. I modified it to arrange for the desired bandwidth (about 8 kHz) and notch after the Nyquist frequency at 36 kHz. I simulated it to get the attached results:

Attachment 1: It shows the input PSD (same as the one posted in the previous elog), the filter transfer function, and The resulting output.

Attachment 2: The circuit schematic. The initial part using THS4131 is a differential LPF and the subsequent RC network is the notch.

Attachment 3: This shows the ratio of the aliased downconverted signal to the the in-band signal, representative of the contamination in each bin. Here too, the aliased signals are negligible as compared to the low frequencies but they are not negligible as compared to the higher frequencies (above 10 kHz) into which they would get downconverted due to sampling. However, here, the attenuation at 8kHz is less than 6 dB while in the previous circuit, it was about 12 dB. One problem with this circuit is at about 6kHz, there is aliased signal from the 65k to 98kHz band, but this can be taken care of by adding an LPF later.

Quote:

Summary: The aim is to design an analog anti-aliasing (AA) filter placed before the ADC, whose function is to filter out components of the input spectrum that have frequencies higher than the Nyquist frequency. This needs to be done so that there is no contamination of aliased downconverted high-frequency signals into the ADC output. I have put down and simulated a circuit to do this, based on the spectra of a few interferometer signals that eric Provided. Attachment 1 shows such an input PSD, treated with whitening filter, before the AA. The sampling rate is 65536 Hz and hence the Nyquist freq. is 32768 Hz.

Motivation: Attachments 2 and 3 show the plot of required attenuation for various frequencies above the Nyquist. We can see a peak at 36 kHz, which will alias to about 29kHz. It will require about 70 dB attenuation here. This indicates that use of a notch filter combined with a low pass filter can be used.

Details of Schematic: Attachment 4 shows the schematic of a Boctor low pass notch filter, cascaded by a 2nd order LPF. The stopband frequency of the boctor filter can be tuned to around 36 kHz. Its main advantage for the boctor is better insensitivity to component value tolerances, use of a single op amp, and relatively independent tuning of parameters.  The various component values are calculated from here. The transfer functions for the circuit shown in attachment 4 were simulated using TINA - a spice based simulation software. The transfer function is shown in attachment 5.

A few more calculations: Attachment 6 shows the output psd after the signal has been treated with AA. Attachments 7 and 8 show the ratio of aliased downconverted signal and the unaliased signal of the output. Here, we can see that above about 13 kHz, the ratios go above -40dB, which is apparently undesirable. However, we also see from the transfer function of the filter that the gain falls to less than -20dB after about this frequency, and the aliased signals are atleast 20 dB lower than this, atleast upto about 29 kHz in attachment 7 and about 25 kHz in attachment 8. This means that the aliased signals are negligible as compared to the low frequencies even if they are not negligible as compared to the higher frequencies (above 13 kHz) into which they would get downconverted due to sampling. But these higher frequencies (above 13 kHz) themselves are small.

The filter overall, is 4th order. Considering this and the above discussion, I need to decide what changes to make in the existing schematic. For now, I could discuss with eric to finalize the opamp and start building the pcb board design.

 

Attachment 1: io.pdf
io.pdf
Attachment 2: AA.JPG
AA.JPG
Attachment 3: ratios_v2.pdf
ratios_v2.pdf
  12211   Wed Jun 22 10:15:45 2016 varunUpdateCDSDAFI update: stereo output

I have updated the DAFI with the following changes:

1) Separated both the channels of stereo output completely, as well as in the GUI.

2) Added text monitors for the inputs and outputs.

The stereo output is now ready except for a cable going from the second channel of the output fibox to the audio mixer.

Attached is the main DAF_OVERVIEW screen and its link button from the LSC screen labelled "DAFI"

Quote:

I wish to have stereo audio output for the DAF module. Hence, there needs to be a second output from the DAF. I added this second output to the model. Following are the details:

FiBox: It consists of two analog inputs which are digitized and multiplexed and transmitted optically. (only 1 fiber is needed due to multiplexing). Attachment 1 shows the fibox with its 2 analog inputs (one of which, is connected), and 1 fiber output. The output of the DAF goes to the FiBox. Until today, the Fibox recieved only 1 analog input. This analog signal comes from the DAC-8 (count starting from 0), which is located at "CH 1 OUT" SMA output in the "MONITORS" bin on the racks (attachment 2).

I have added another output channel to the DAF model both in software and in hardware. The DAF now also uses DAC-9 analog output which goes to the second analog input of the FiBox. The DAC-9 output is located at "CH 2 OUT" SMA output in the "MONITORS" bin on the racks (attachment 4).

After making the changes, the Fibox is shown in attacment 3.

Testing: The LSC input on passing through the DAF block is given through two different DAC outputs, to the same Fibox channel (one after the other), and the output is heard. More concrete testing will be done tomorrow. It will be as follows:

1) Currently, I need to search for a suitable cable that would connect the second channel of the output fibox to the audio mixer. After doing this, end to end testing of both channels will be done.

2) I could not access the AWG, probably because the DAQ was offline today afternoon. Using a signal from the AWG will give a more concrete testing of the stereo output.

3) After this, I will separate the two channels of the stereo completely (currectly they are seperated only at the DAF output stage)

4) I also will edit the medm gui appropriately.

 

Quote:

I have added Enable buttons for each of the DSP blocks, and labels for the matrix elements. The input matrix takes inputs from each of the 4 channels: ADC1, ADC2, LSC and EXC, and routes them to the audio processing blocks (attachment 2). The output matrix (attachment 3) takes the outputs of the various DSP blocks and routes them to the output and then to the speakers. 

 

 

Attachment 1: C1DAF_OVERVIEW.png
C1DAF_OVERVIEW.png
Attachment 2: DAF_link_from_LSC.png
DAF_link_from_LSC.png
  12210   Wed Jun 22 08:40:42 2016 ranaUpdateLSCY arm @ 30kHz UGF w/POY, AO

Below 100 Hz, I suppose this means that the X arm is now limited by the quadrature sum of the X and Y arm seismic noise.

  12209   Tue Jun 21 14:12:10 2016 SteveUpdatePEM vacuum envelope wiped

I wiped down the cranes with wet towel and Mario our janitor did the chamber tops with the tubes.

The optical tabels were not touched.

The outside temp peaked at 44 C yesterday

 

Attachment 1: 60dPEM.png
60dPEM.png
  12208   Tue Jun 21 11:49:29 2016 ericqFrogsCDSmedm command not working

The workstations' .bashrc is a symbolic link to /users/controls/.bashrc

In it, someone commented out the critical line:

#source /ligo/cdscfg/workstationrc.sh

I uncommented it. medm (and all of the other things like cdsutils) work again.

I blame jamie.

  12207   Tue Jun 21 11:26:42 2016 varunFrogsCDSmedm command not working

"medm: command not found" error when run through command line both in pianosa and rossa in both editing and execution modes. It however gets executed and edited through the sitemap button. Don't know the source of the problem. Gautam did check the .bashrc file. aliases for SITEMAP and m40m are intact in the .bashrc file.

ELOG V3.1.3-